Base bibliographique en relation avec les bolomètres pour la mission Planck HFI

A. Coulais

May 31, 2005

Cette liste volontairement limitée aux articles majeures a été élaborée en relation avec M. Piat. N’hésitez pas à nous contacter pour tout échange sur ces sujets.

References

ChaninTorre84 [CT84] G. Chanin and J. P. Torre. Electrothermal model for ideal semiconductor bolometers. Optical Society of America Journal, 1:412–419, April 1984.

1990fgbs.rept..359G [GAC^{+}90] M. Griffin, P. Ade, S. Church, A. Murray, M. Overhamm, R. Faymonville, and S. Wang. Performance tests on Ge:Ga and Ge:Be detectors for the ISO long wavelength spectrometer. In From Ground-Based to Space-Borne Sub-mm Astronomy, pages 359–364, December 1990.

1981ApOpt..20.3992M [Mat81] J. C. Mather. Bolometer cavity design for efficiency. Applied Optics, 20:3992–3993, December 1981.

1982ApOpt..21.1125M [Mat82] J. C. Mather. Bolometer noise: nonequilibrium thoery. Applied Optics, 21:1125–1129, March 1982.

1984ApOpt..23..584M [Mat84] J. C. Mather. Bolometers: ultimate sensitivity, optimization, and amplifier coupling. Applied Optics, 23:584–588, February 1984.

1978acge...23..640T [TC78] J. P. Torre and G. Chanin. Test flight results of a balloon-borne He-3 cryostat. In Advances in cryogenic engineering. Volume 23 - Proceedings of the Conference, Boulder, Colo., August 2-5, 1977. (A79-15301 04-31) New York, Plenum Press, 1978, p. 640-643., pages 640–643, 1978.

1984RScI...55..213T [TC84] J. P. Torre and G. Chanin. Heat switch for liquid-helium temperatures. Review of Scientific Instruments, 55:213–215, February 1984.

1985RScI...56..318T [TC85] J. P. Torre and G. Chanin. Miniature liquid-^{3}He refrigerator. Review of Scientific Instruments, 56:318–320, February 1985.

Woodcraft03 [WSA^{+}03] A. L. Woodcraft, R. V. Sudiwala, P. A. R. Ade, M. J. Griffin, E. Wakui, R. S. Bhatia, A. E. Lange, J. J. Bock, A. D. Turner, M. H. Yun, and J. W. Beeman. Predicting the response of a submillimeter bolometer to cosmic rays. Applied Optics, 42:5009–5016, September 2003.